You sprained your ankle and need to go to the Emergency Room and it will cost $10,000 without insurance. Your health insurance has a $1,000 deductible, emergency room co-pay of $750, and your co-insurance is 70/30. How much of the $10,000 will you be responsible for?
Answer #1

Please give a thumbs up if you find this helpful :)

Correct Answer: $ 4,750

Working :

Parameters provided in the question,

Cost without insurances = $ 10,000

Deductible = $ 1,000

Emergency co-pay = $ 750

Co-insurance = 70/30

Brief summary of different pays:

  • A deductible is an amount paid for health care services before your health insurance begins to pay
  • Co-Insurance of 70/30 depicts that your plan will cover 70 percent cost of the bills and other 30 percent will be borne by you
  • Copay is a fixed amount you pay for the service when you avail them

The amount responsible by you = Deductible + Co-insurance (30%) + Emergency room co-pay

The amount responsible by you = 1,000 + 30 % of 10,000 + 750

The amount responsible by you = 1,000 + 3,000 + 750

The amount responsible by you = $ 4,750

Thus amount you will be responsible including all the deductible will be $ 4,750